You determine that you have only three copies left of an important DNA fragment, so you decide to amplify it. Using flanking primers, how many PCR cycles would you have to run to generate over one billion (109) copies of the fragment? I know the answer is 29. I just don't understand how to get that answer.
Answer #1
  • During each cycle, the DNA gets doubled.
  • So, after one cycle it will be 2 x 2 =4 ( if the cycle is exponential)
  • After 10 cycles, it is 210 = 1024
  • After 20 cycles, it is 220 = 1048576
  • Number of DNA copies produced after a prescribed number of cycles, if it is exponential = 2(n+1)
  • Where n is number of cycles and one is DNA original fragment already present.
  • So, after 29 cycles = 2(29+1) = 2 (30) = more than 1 billion
  • 2n = 230 = 1.02 x 109
  • But here we have started with 3 molecules of DNA fragments, so it is 3 times 2n or ...... 3 x 2(n+1)
  • Then, after 10 cycles = 3 x 2(10) = 3072
  • after 20 cycles, 3 x 220=3145728
  • after 25 cycles , 3x 213 = 1.0 x 109
  • So, the correct answer is 25 cycles not 29.